OpenStudy (anonymous):

Solve using the elimination method 3r-4s=-6 4r+3s=17 This would not be possible right? or would it be 7r-1s=9?

OpenStudy (anonymous):

eliminate either r or s...s looks better since you have opposite signs already

OpenStudy (anonymous):

multiply by top by 3 and bottom by 4

OpenStudy (anonymous):

would it be r=1.3?

OpenStudy (anonymous):

9r - 12s = -18 16r +12s=68

OpenStudy (anonymous):

solve for r

OpenStudy (anonymous):

r= 10/3 sp sub that in to find s

OpenStudy (anonymous):

25r=50 r=2?

OpenStudy (anonymous):

3(10/3) - 4s = -6 10-4s=-6 -4s = -16 s= 4

OpenStudy (anonymous):

(10/3, 4)

OpenStudy (anonymous):

Ok I'm really confused, where did the 10 come from? Sorry I have a learning disability and really have a hard time with math.

OpenStudy (anonymous):

it is solving a system by elimination with multiplication....ever heard of that

OpenStudy (anonymous):

Yes, I am doing that right now in my math class, please don't think I am looking for you to give me the answer. I have Math 117 next and I have to understand this.

OpenStudy (anonymous):

It has only ever shown how to solve when y or x is a negative and positive of the same number never two different numbers

OpenStudy (anonymous):

can you come to whiteboard so I Can write it out http://www.twiddla.com/536745

OpenStudy (anonymous):

(2,3)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

YW!!!